ch 24 Q 1.) The team monitoring a space probe exploring the outer solar system finds that radio trasmissions from the probe take 2.62 hours to reach earth. How distant (in meters) is the probe?
Answer #1

The elapsed time t is given in hours, so let us first convert into seconds

t = 2.62*3600 = 9432 sec

The distance between the earth an the probe is

d = c*t

= 3*10^8*9432

= 2.83*10^12 m